40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon May 9 15:21:36 2011, josephb, Update, PSL, Fixed channel names for PSL QPDs, fixed saturation, changed signs 
    Reply  Tue May 10 10:06:23 2011, josephb, Update, PSL, Fixed channel names for PSL QPDs, fixed saturation, changed signs 
Message ID: 4666     Entry time: Mon May 9 15:21:36 2011     Reply to this: 4677
Author: josephb 
Type: Update 
Category: PSL 
Subject: Fixed channel names for PSL QPDs, fixed saturation, changed signs 

[Valera, Joe]

Software Changes:

First we changed all the C1:IOO-QPD_*_* channels to C1:PSL-QPD_*_* channels in the /cvs/cds/caltech/target/c1iool0/c1ioo.db file, as well as the /opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini file.  We then rebooted the frame builder via "telnet fb 8087" and then "shutdown".

This change breaks continuity for these channels prior to today.

The C1:PSL-QPD_POS_HOR and C1:PSL-QPD_POS_VERT channels were found to be backwards as well.  So we modified the /cvs/cds/caltech/target/c1iool0/c1ioo.db file to switch them.

Lastly, we changed the ASLO and AOFF values for the C1:PSL-QPD_POS_SUM and the C1:PSL-QPD_ANG_SUM so as to provide positive numbers.  This was done by flipping the sign for each entry.

ASLO went from 0.004883 to -0.004883, and AOFF when from -10 to 10 for both channels.

Hardware Changes:

The C1:PSL-QPD_ANG_SUM channel had been saturated at -10V.  Valera reduced the power on the QPD to drop it to about 4V by placing an ND attenuator in the ANG QPD path.

ELOG V3.1.3-